I'm helping my son solve his inequalities assignment, I'm not sure of the answer can you please verify it. thanks so much. The problem is:
4(3x+5)-3>= 6(2x+3)-1
12x+20-3>=12x+18-1
12x+17>=12x+17
x>=0

Answer by Juhi(44)

4(3x+5)-3> = 6(2x+3)-1
12x+20-3> = 12x+18-1
12x+17> = 12x+17
x> = 0